# Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation

### v1

@@ 4.9 Case Study
We use two cases that are beneficial for us in understanding our methods’ limitations. See more descriptions in Appendix H.
@@ 5 Conclusion
Given the limitations of previous methods in accurately retrieving knowledge triples and their challenges in handling dialogue tasks due to coreference issues, we propose a novel framework. This framework features a carefully designed knowledge retriever to identify valuable knowledge triples, an advanced dialogue rewriter to resolve coreference effectively, and a knowledge-enhanced dialogue generation component to improve factuality.
Building on this, we also adapt the fact score to make it suitable for evaluating the factuality of the dialogue response and verifying its validity in the dialogue task.
The results demonstrate that our method significantly enhances the factual accuracy of dialogue responses compared to other baselines, including the SOTA G-retriever.
We analyse some cases and see some limitations in our current work. We plan to explore the mechanism of LLM reactions to triples in the future.
Limitations
Although we propose a well-designed framework for dialogue-related tasks, some limitations remain.
Our current knowledge retrieval method relies on the exact entity overlap between the query and Wikidata. This approach lacks semantic understanding, making it inefficient in retrieving semantically related knowledge triples. Constructing a graph database from Wikidata could be an effective way to address this problem, but building such a large-scale database presents significant challenges.
Additionally, our approach relies on knowledge triples. However, the validity of certain triples, such as (“Montevideo”, “population”, “309331”) shown in the key knowledge of Case 1 in Section 4.9, will change over time. Retrieving such triples for LLMs could lead to outdated responses.
Ethical Statement
In this work, we follow the ethical standards in data use and research. The datasets used in this study, OpendialKG and HybriDialogue, are publicly available and have been widely used in research. We have not added any extra personal information to these publicly available datasets.
The data from human annotators was limited to some specific labels, and the results reported in this paper are based on statistical analysis. No personal or sensitive information was accessed during the research process.
@@ Appendix A Prompts
The prompts of the knowledge-augmented approach are shown in Table 6. The factuality-related prompts are described in Table 8.
Dialogue Rewrite Prompt
@@
Table 6: Prompt for rewriting the dialogue and generating a response with dialogue and knowledge triples.
@@
Knowledge: {Selected Triples} Dialogue: {Dialogue Context} Given the above knowledge and dialogue, please respond to the input below and ensure the response is fluent and fact-consistent in English. Input: {Query} Response: … {Examples..} If the following input is an incomplete sentence or a phrase, please output it exactly as it is. Otherwise, if it is a complete sentence, split it into atomic sentences based only on the given information, without adding any additional information or making inferences: Input: {Response} Output …:
Table 7: The prompts for getting atomic facts and fact score.
Fact Score Prompt
Instruction: The statement is part of a response in a dialogue. Evaluate the statement strictly based on the provided knowledge source and dialogue history only. If the statement is not a factual claim (e.g., opinion, question, or unclear assertion), output: "no enough information." If it is a factual claim: Output true if the statement is directly supported by evidence in the knowledge source or dialogue history. Output false if the statement is directly contradicted by the knowledge source or dialogue history. Output no enough information if there is no direct evidence for or against the statement. Important: Do not use your intern knowledge or make inferences. Please only output your final answer and do not output any explanations. Evidence: {Wikipedia Passages} Dialogue history: {Dialogue} Speaker A: {Speaker} Statement: {Atomic Fact}
Table 8: The prompts for getting atomic facts and fact score.
@@ Appendix B Datasets Details
It is an English knowledge-driven dialogue dataset that consists of a recommendation task related to movies and books and a chit-chat task related to sports and music. The dataset contains 13,802 samples. After processing, some samples are not valid in the task, so we remove them. Finally, we randomly selected 15% (1,962 samples) for the validation set, another 15% (1,973 samples) for the test set, and the remaining 70% (9,120 samples) for the training set. The dataset provides a number of triples extracted from Freebase Bast et al. (2014). These triples were collected several years ago. To address any potential issues with outdated information, we re-extract the triples from Wikidata. OpendialKG is released under CC-BY-NC-4.0 license [^2] https://github.com/facebookresearch/opendialkg, which permits non-commercial research use.
@@
It is an open-domain and information-seeking dialogue dataset in English. It is constructed by splitting complex questions into multi-turn dialogue. The original dataset does not offer triples, so we collect them by matching the entities with triples from Wikidata. The training set contains 4,359 samples, while the validation set includes 242 samples, and the test set consists of 243 samples. HybriDial follows the MIT license Nakamura et al. (2022), which allows both commercial and non-commercial use.
@@
Query: Could you recommend some movies staring Chiaki Kuriyama?
Retrieved triples from BM25: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, family name, Kuriyama) (Chiaki Kuriyama, place of birth, Tsuchiura) (Chiaki Kuriyama, occupation, film actor) (Chiaki Kuriyama, instrument, voice) …
Retrieved triples from KAPING: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, occupation, model) (Chiaki Kuriyama, occupation, singer) (Chiaki Kuriyama, occupation, actor) (Chiaki Kuriyama, occupation, fashion model) …
Retrieved triples from Triton: (Kill Bill Volume 1, cast member, Chiaki Kuriyama), (Into the Sun, cast member, Chiaki Kuriyama), (Kagen no Tsuki, cast member, Chiaki Kuriyama), (Kamogawa Horumo, cast member, Chiaki Kuriyama), (Gonin, cast member, Chiaki Kuriyama) …
@@
Query: Hi. Can you tell me who Judd Trump is?
Retrieved triples from BM25: (Judd Trump, given name, Judd)(Judd Trump, family name, Trump)(2021 Champion of Champions, winner, Judd Trump)(Judd Trump, victory, 2016 China Open)(Judd Trump, victory, 2011 China Open) …
Retrieved triples from KAPING: (Judd Trump, occupation, snooker player) (Judd Trump, nickname, The Ace) (Judd Trump, given name, Judd) (Judd Trump, family name, Trump) (Judd Trump, place of birth, Whitchurch) …
Retrieved triples from Triton: (2019 World Snooker Championship, winner, Judd Trump), (Judd Trump, award received, Snooker Hall of Fame), (Judd Trump, country of citizenship, United Kingdom), (Judd Trump, sport, snooker), (Judd Trump, award received, player of the year award) …
@@
Appendix C Examples from GPT-4o
We employ GPT-4o to annotate query-triple pairs and resolve dialogue coreference from the training dataset. In the end, we collected 29K samples of query-triple pairs and 35.4K samples of coreference resolution with reasoning steps. The examples for fine-tuning are listed in Tables 10 and  11.
Query Triple Relevance
@@ Appendix C Examples from GPT-4o
Table 10: The query and triple are sourced from the training dataset and Wikidata. The relevance is annotated by GPT-4o.
@@
CoT Speaker A mentions "Lev Grossman" as the author they like. This establishes that "his" in their request for book recommendations refers to Lev Grossman. Speaker B mentions "He wrote The Magician’s Land and The Magicians." The pronoun "He" refers to Lev Grossman, as he is the author being discussed. Speaker A uses "it" in "I think I’ve read The Magicians it a High Fantasy book." The pronoun "it" refers to "The Magicians," as it is the last explicitly mentioned book in the conversation.
Resolved Dialogue Speaker A: I like Lev Grossman books can you recommend some of Lev Grossman’s books Speaker B: Sure! Lev Grossman wrote The Magician’s Land and The Magicians. Speaker A: Oh, I think I’ve read The Magicians. The Magicians is a High Fantasy book.
Table 11: An example generated by GPT-4o illustrates how CoT reasoning works in coreference resolution.
Appendix D Experimental Setup
For fine-tuning the Triton, we only collect the query-triple pairs from the HybriDial dataset. We fine-tuned 3 epochs for either the LoRA method or the full-tuning method. The rank and alpha parameters for LoRA are both set to 8.
We utilise the multi-qa-mpnet-base-dot-v1 version for the baseline KAPING. As G-retriever is implemented using Llama2 7B Touvron et al. (2023), we adopt the same model in our Triton-X under identical settings to ensure fairness. For knowledge-augmented methods, the maximum number of selected triples is set to 5. Additionally, all experiments were run a single time.
Regarding the evaluation metrics, we utilised the evaluate package  [^3] https://huggingface.co/docs/evaluate to compute BLEU-4 and ROUGE-L scores. Besides, the PPL is calculated using GPT-2 Radford et al. (2019).
We fine-tune our relevance classifier using a single A100 GPU and our dialogue generation model with three A100 GPUs. We utilized two A100 GPUs for inference tasks. All A100 GPUs we used in this work have 80GB of memory. In total, our GPU runtime amounted to approximately 575 hours.
Appendix E Conventional Evaluation Metrics in Detail
@@
  • BLEU Reiter (2018), used in the machine translation initially, calculates the n-gram overlap between generated text and references and reflects the correlation between generated text and human writing.
@@
  • ROUGE-L Lin (2004) is calculated based on the length of the longest common subsequence, which captures the word order and sentence-level structure from the ground truth.
@@
  • PPL Jelinek et al. (1977) is widely used in measuring text fluency of generated output, in which calculation is based on the language model.
@@
  • F1 Score Nan et al. (2021) is based on the precision and recall of the extracted entities matching between the generated response and ground truth.
Case 1 Case 2
Context: A: What is the capital and largest city in Uruguay? B: Montevideo is the capital and largest city in Uruguay. Query: What was the population of Montevideo, Uruguay, in 2011? Response: The population of Montevideo, Uruguay, in 2011 was 59,027. Key Knowledge: (Montevideo, population, 309331) New Response: Montevideo, population, 309331 Dialogue: A: What kind of game was “Team Fortress 2”? B: It’s a multiplayer first-person shooter. A: And what was it’s release date? Response: It was released on October 26, 2007. Rewritten Dialogue: B: What kind of game was “Team Fortress 2”? A: “Team Fortress 2” is a multiplayer first-person shooter. A: And what was “Team Fortress 2”’s release date? New Response: “Team Fortress 2” was released on November 20, 2007.
Table 12: The responses of cases were both generated by Flan-T5-Large. The new response was created using key knowledge or a rewritten dialogue. Triton selected the key knowledge.
Appendix F Details of Manual Annotation
@@
We invited some students from Asian countries to annotate samples. All of them are well-educated and possess good English proficiency. The annotators were told that the annotated data would be used for research, and we got consent for data use.
@@
To assess the accuracy of query-triple pair annotated from GPT-4o, we invite tree annotators to this task. We randomly selected 100 samples and instructed participants to determine whether the query and triple were relevant. The output is either relevant or irrelevant. In the beginning, two annotators have 0.83 in agreement and 0.66 in Cohen’s Kappa. We introduced another annotator to mediate the disagreement. We utilized the final annotated result to assess the accuracy of query-triple pairs, which is 0.81 in agreement and 0.62 in Cohen’s Kappa. These results indicate a high level of accuracy.
@@
Three annotators were involved in the assessment of the fact score task. We used the fact score prompt, described in Table 8, as an instruction. Two annotators independently assessed the fact score for 100 samples from the HybriDial and OpendialKG datasets respectively. Initially, the raw agreement scores were 0.76 for HybriDial and 0.81 for OpenDialKG. Their inter-annotator agreement, measured by Cohen’s Kappa, was 0.613 for OpendialKG and 0.705 for HybriDial. To resolve discrepancies, we introduced a third annotator to mediate disagreements. Finally, we report the final agreement and Cohen’s Kappa score with the evaluation model, as shown in Table 1.
Two annotators were involved in the dialogue human evaluation. The accuracy of human annotations is consistently high across all aspects, with HybriDial achieving 0.81 for coherence, 0.88 for fluency, and 0.89 for informativeness, while OpendialKG reaches 0.83, 0.96, and 0.80 respectively, indicating the overall reliability of manual evaluations. We finally report the average score in our dialogue human evaluation.
Appendix G Baseline Descriptions

### v2

@@ 4.9 Case Study
We use a case about Hayden Panettiere shown in Table 5 to illustrate our work. From the case, the dialogue contains pronouns, e.g. “she”, challenging the generation of factually correct responses. We can see the response from Flan-T5-XXL is repeating the previous utterance “She played a voice on the cartoon A Bug’s Life…”. After dialogue reformulation and dialogue sense knowledge selection, our TG-DRG-R is able to generate a more factual and detailed response. However, from Table 2, we observe that TG-DRG-R’s response achieves lower BLEU-4 and ROUGE-L scores compared to those of Flan-T5-XXL. This suggests that BLEU-4 and ROUGE-L may not effectively capture the quality of responses in open-domain dialogue tasks.
@@ 5 Conclusion
Our work addresses the limitations of previous methods in handling factual dialogue response generation, particularly issues arising from complex coreference in the dialogue. To this end, we propose two novel frameworks, TG-DRG and GA-DRG, which feature reasoning-guided dialogue reformulation and dialogue sense knowledge selection, leveraging graph-based generation techniques. The results demonstrate that these frameworks noticeably enhance the factual accuracy of dialogue responses compared to baseline methods, including the SOTA G-retriever.
Additionally, we propose the dialogue fact score as a novel metric to evaluate the factuality of dialogue responses. This metric shows high consistency with human evaluations and proves to be more reliable than traditional metrics such as BLEU-4, ROUGE-L, and F1, which often fail to capture factual consistency in open-domain dialogue tasks.
While our approach shows noticeable improvements, certain limitations remain. For instance, the validity of triples from external knowledge graphs may change over time, potentially impacting factuality. Addressing this challenge will be a focus of our future work.
@@ Appendix A Prompts
In this section, we show the prompts used in our work.
The prompt for reasoning-guided dialogue reformulation is described in Table 6. The prompts for the knowledge-augmented dialogue response generation, used in TG-DRG, and atomic fact splitting are shown in Table 7. The prompt for fact verification of dialogue fact score prompts is described in Table 8.
Dialogue Reformulation Prompt
@@
Table 6: The prompt for reasoning-guided dialogue reformulation.
@@
Knowledge: {Selected Triples} Dialogue: {Dialogue Context} Given the above knowledge and dialogue, please respond to the input below and ensure the response is fluent and fact-consistent in English. Input: {Last Utterance} Response: … {Examples..} If the following input is an incomplete sentence or a phrase, please output it exactly as it is. Otherwise, if it is a complete sentence, split it into atomic sentences based only on the given information, without adding any additional information or making inferences: Input: {Response} Output …:
Table 7: The prompts for dialogue response generation (left) and atomic fact splitting (right).
Dialogue Fact Score Prompt
Instruction: The statement is part of a response in a dialogue. Evaluate the statement strictly based on the provided knowledge source and dialogue history only. If the statement is not a factual claim (e.g., opinion, question, or unclear assertion), output: ”no enough information.” If it is a factual claim: Output true if the statement is directly supported by evidence in the knowledge source or dialogue history. Output false if the statement is directly contradicted by the knowledge source or dialogue history. Output no enough information if there is no direct evidence for or against the statement. Important: Do not use your intern knowledge or make inferences. Please only output your final answer and do not output any explanations. Evidence: {Wikipedia Passages} Dialogue history: {Dialogue} Speaker A: {Speaker} Statement: {Atomic Fact}
Table 8: The prompt for fact verification of dialogue fact score.
@@ Appendix B Datasets Details
It is a knowledge-driven dialogue dataset. The dataset contains 13,802 samples. After processing, some samples are not valid in the task, so we remove them. Finally, we randomly selected 15% (1,962 samples) for the validation set, another 15% (1,973 samples) for the test set, and the remaining 70% (9,120 samples) for the training set. The dataset provides a number of triples extracted from Freebase (Bast et al. 2014). These triples were collected several years ago. To address any potential issues with outdated information, we re-extract the triples from Wikidata. OpendialKG is released under CC-BY-NC-4.0 license, which permits non-commercial research use.
@@
It is an open-domain dialogue dataset. It is constructed by splitting complex questions into multi-turn dialogue. The original dataset does not offer triples, so we collect them by matching the entities with triples from Wikidata. The training set contains 4,359 samples, while the validation set includes 242 samples, and the test set consists of 243 samples. HybriDial follows the MIT license (Nakamura et al. 2022), which allows both commercial and non-commercial use.
@@
Last Utterance: Could you recommend some movies staring Chiaki Kuriyama?
Selected triples from BM25: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, family name, Kuriyama) (Chiaki Kuriyama, place of birth, Tsuchiura) (Chiaki Kuriyama, occupation, film actor) (Chiaki Kuriyama, instrument, voice) …
Selected triples from KAPING: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, occupation, model) (Chiaki Kuriyama, occupation, singer) (Chiaki Kuriyama, occupation, actor) (Chiaki Kuriyama, occupation, fashion model) …
Selected triples from our method: (Kill Bill Volume 1, cast member, Chiaki Kuriyama), (Into the Sun, cast member, Chiaki Kuriyama), (Kagen no Tsuki, cast member, Chiaki Kuriyama), (Kamogawa Horumo, cast member, Chiaki Kuriyama), (Gonin, cast member, Chiaki Kuriyama) …
@@
Last Utterance: Hi. Can you tell me who Judd Trump is?
Selected triples from BM25: (Judd Trump, given name, Judd)(Judd Trump, family name, Trump)(2021 Champion of Champions, winner, Judd Trump)(Judd Trump, victory, 2016 China Open)(Judd Trump, victory, 2011 China Open) …
Selected triples from KAPING: (Judd Trump, occupation, snooker player) (Judd Trump, nickname, The Ace) (Judd Trump, given name, Judd) (Judd Trump, family name, Trump) (Judd Trump, place of birth, Whitchurch) …
Selected triples from our method: (2019 World Snooker Championship, winner, Judd Trump), (Judd Trump, award received, Snooker Hall of Fame), (Judd Trump, country of citizenship, United Kingdom), (Judd Trump, sport, snooker), (Judd Trump, award received, player of the year award) …
@@
Appendix C Examples generated from GPT-4o
We employ GPT-4o to generate data for dialogue sense knowledge selection and dialogue reformulation from the training dataset, where we collected 29K samples and 35.4K for them. The examples for fine-tuning are listed in Tables 10 and  11.
Last Utterance Triple Relevance
@@ Appendix C Examples generated from GPT-4o
Table 10: The dialogue and corresponding triple are taken from the training dataset and Wikidata, respectively, with only the last utterance of the dialogue used here.. The relevance is generated by GPT-4o. These examples are adopted for fine-tuning our dialogue sense knowledge selection.
@@
CoT Speaker A mentions ”Lev Grossman” as the author they like. This establishes that ”his” in their request for book recommendations refers to Lev Grossman. Speaker B mentions ”He wrote The Magician’s Land and The Magicians.” The pronoun ”He” refers to Lev Grossman, as he is the author being discussed. Speaker A uses ”it” in ”I think I’ve read The Magicians it a High Fantasy book.” The pronoun ”it” refers to ”The Magicians,” as it is the last explicitly mentioned book in the conversation.
Reformulated Dialogue Speaker A: I like Lev Grossman books can you recommend some of Lev Grossman’s books Speaker B: Sure! Lev Grossman wrote The Magician’s Land and The Magicians. Speaker A: Oh, I think I’ve read The Magicians. The Magicians is a High Fantasy book.
Table 11: An example generated by GPT-4o illustrates how CoT reasoning works in dialogue reformulation.
Example 1
Dialogue:
Speaker A: I enjoy Charlaine Harris’s work, can you name something similar?
Speaker B: The Southern Vampire Mysteries I would recommend for you.
Speaker A: When was it released?
Speaker B: 2004.
Speaker A / Reformulated
How many books are in the series? / How many books are in The Southern Vampire Mysteries series?
Label: Good
Reason: The reference “The series” in the last utterances has been resolved as “The Southern Vampire Mysteries” correctly.
Example 2
Speaker A: Could you recommend a book written by author Fannie Flagg?
Speaker B: Welcome to the World, Baby Girl and I Still Dream About You are great books that I would suggest. You read them?
Speaker A: No, I have read neither of those. What are their genres?
Speaker B: It is adventure and is also a movie released in 2010, I would suggest reading the book then watching the movie.
Speaker A / Reformulated:
Yes, I will definitely read the book first. Can you add it to my list? / Yes, I will definitely read “Welcome to the World, Baby Girl” first. Can you add “Welcome to the World, Baby Girl” to my list?
Label: Neutral
Reason: The reference “the book” in the dialogue is ambiguous. So the reformulated utterance may not be correct.
Example 3
Dialogue:
Speaker A: Cool! do you know the occasion name of the Video event?
Speaker B: Yes, it is NeXT’s internal marketing strategy video.
Speaker A: Great! do you know the city where that video took place?
Speaker B: Yes, it was Redwood City.
Speaker A / Reformulated:
Good! do you know how many miles that city far from south of San Francisco? / Good! do you know how many miles that city far from south of San Francisco?
Label: Bad
Reason: The reference “that city” should be resolved, but the model fails to resolve it.
Table 12: The manual annotated examples for dialogue reformulation. In human evaluation, we only evaluate the last utterance.
Appendix D Conventional Evaluation Metrics in Detail
@@
  • BLEU (Reiter 2018), used in the machine translation initially, calculates the n-gram overlap between generated text and references and reflects the correlation between generated text and human writing.
@@
  • ROUGE-L (Lin 2004) is calculated based on the length of the longest common subsequence, which captures the word order and sentence-level structure from the ground truth.
@@
  • PPL (Jelinek et al. 1977) is widely used in measuring text fluency of generated output, in which calculation is based on the language model.
@@
  • F1 Score (Nan et al. 2021) is based on the precision and recall of the extracted entities matching between the generated response and ground truth.
Case 2
Original Dialogue:
Speaker A: Hello! do you know fictional marsupials?
Speaker B: Yes, it is subsidiary to the list of fictional animals and is a collection of various notable marsupial characters that appear in various works of fiction.
Speaker A: Great! can you list some literature names of fictional marsupials?
Speaker B: Yes, there are many books available including Ginger Kangaroo, Kidding Kangaroo, etc.
Speaker A: Good! do you know the species name of the literature Blinky Bill?
Speaker B: Yes, it is Koala.
Speaker A: Great! do you know the author of that book?
Flan-T5-XXL: Yes, it is written by E. Nesbit.
Fact Verification: False
Reformulated Dialogue:
Speaker A: Hello! do you know fictional marsupials?
Speaker B: Yes, fictional marsupials are subsidiary to the list of fictional animals and is a collection of various notable marsupial characters that appear in various works of fiction.
Speaker A: Great! can you list some literature names of fictional marsupials?
Speaker B: Yes, there are many books available including Ginger Kangaroo, Kidding Kangaroo, etc.
Speaker A: Good! do you know the species name of the literature Blinky Bill?
Speaker B: Yes, the species name of Blinky Bill is Koala.
Speaker A: Great! do you know the author of Blinky Bill?
Key Knowledge: (Blinky Bill, illustrator, Dorothy Wall), (Blinky Bill, author, Dorothy Wall) …
TG-DRG-R: Yes, the author of Blinky Bill is Dorothy Wall.
Fact Verification: True
Case 3
Context:
Speaker A: What is the capital and largest city in Uruguay?
Speaker B: Montevideo is the capital and largest city in Uruguay.
Speaker A: What was the population of Montevideo, Uruguay, in 2011?
Flan-T5-Large: The population of Montevideo, Uruguay, in 2011 was 59,027.
Key Knowledge: (Montevideo, population, 309331)
TG-DRG-NR: Montevideo, population, 309331
Table 13: The responses of cases were generated by Flan-T5 series models. In case 2, the resolved coreference is bold in the reformulated dialogue.
Appendix E Details of Manual Annotation
@@
We invited several annotators to annotate the samples. All of them are well-educated and possess good English proficiency. The annotators were told that the annotated data would be used for research, and we got consent for data use.
@@
To assess the accuracy of the samples used in fine-tuning dialogue sense knowledge selection generated from GPT-4o, we invite three annotators to this task. We randomly selected 100 samples and instructed participants to determine whether the last utterance of the dialogue is relevant to the triple. The output is either relevant or irrelevant. In the beginning, two annotators have 0.83 in agreement and 0.66 in Cohen’s Kappa. We introduced another annotator to mediate the disagreement. We utilized the final annotated result to assess the accuracy of query-triple pairs, which is 0.81 in agreement and 0.62 in Cohen’s Kappa. These results indicate a high level of accuracy.
@@
Three annotators were involved in the assessment of the dialogue fact score task. We used the dialogue fact score prompt, described in Table 8, as an instruction. Two annotators independently assessed the dialogue fact score for 100 samples from the HybriDial and OpendialKG datasets, respectively. Initially, the raw agreement scores were 0.76 for HybriDial and 0.81 for OpenDialKG. Their inter-annotator agreement, measured by Cohen’s Kappa, was 0.613 for OpendialKG and 0.705 for HybriDial. To resolve discrepancies, we introduced a third annotator to mediate disagreements. Finally, we report the final agreement and Cohen’s Kappa score with the evaluation model, as shown in Table 1.
We randomly selected 100 samples to evaluate the quality of dialogue reformulation, with two annotators involved in the process. The evaluation focuses solely on the last utterance, as it is the most crucial part of the dialogue. Each sample is manually rated as good, neutral, or bad. On the HybriDialogue dataset, the annotators achieved a raw agreement of 0.91 and a Cohen’s kappa of 0.580. On the OpenDialKG dataset, they reached a raw agreement of 0.94 and a Cohen’s kappa of 0.548, indicating moderate agreement. The annotation examples of dialogue reformulation are shown in Table 12.
Appendix F Baseline Descriptions